Circular No. 7016

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998eb IN NGC 1961 W. D. Li, M. Modjaz, E. Halderson, T. Shefler, J. Y. King, M. Pa-

penkova, R. R. Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova in NGC 1961 during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998eb was found and confirmed on unfiltered images taken on Sept. 17.5 and 19.4 UT (unfiltered magnitude ∼ 17.8), respectively. The object, is located at α = 5^{h}42^{m}12^{s}.02, δ = +69^{o}22′26′′.3 (equinox 2000.0), which is ∼ 38′′.8 east and 17′′.0 south of the nucleus of NGC 1961. A KAIT image of NGC 1961 obtained on Mar. 9.2 shows nothing at the position of SN 1998eb (limiting mag ∼ 18.5).

XTE J1946+274 AND 3A 1942+274 T. Takeshima, NASA/Goddard Space Flight Center and Universities

Space Research Association; and D. Chakrabarty, Massachusetts Institute of Technology, report: “Scans of the region containing this transient x-ray pulsar (IAUC 7014) with the Proportional Counter Array (PCA) onboard RXTE on Sept. 16.23 UT have yielded the following improved position for the x-ray source: α = 19^{h}45^{m}34^{s}, δ = +27^{o}23′.0 (equinox 2000.0). The error radius of 2′.4 (90-percent confidence level) is dominated by systematic uncertainties due to source intensity variations during the scans. The error circle almost includes the overlap of the RXTE/ASM error box for XTE J1946+274 (IAUC 7014) and the long Ariel V error strip for 3A 1942+274 (Warwick et al. 1981, MNRAS 197, 865). This strongly suggests that the two sources are identical. During the PCA observations on Sept. 16.24 and 17.57, the average source flux level and the pulse fraction stayed at ∼ 110 mCrab and ∼ 30 percent, respectively (2–60 keV). Observations at other wavelengths are encouraged in order to identify the mass donor.”

Corrigendum. On IAUC 7014, line 30, for 2.2(2) × 10^{−12} Hz/s read 8.7(9) × 10^{−12} Hz/s

COMET C/1997 J2 (MEUNIER-DUPOUY) Total visual magnitude estimates: Aug. 3.29, 11.0 (C. E. Spratt, Vic-

toria, BC, 0.20-m reflector); 19.87, 11.3 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); Sept. 9.81, 11.2 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 18.04, 11.8 (J. Bortle, Stormville, NY, 0.41-m reflector).

1998 September 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green